# Best of both worlds: Fusing hyperspectral data from two generations of spectro-imagers for X-ray astrophysics

### v1

@@ 5 Conclusions
We have developed an algorithm for fusing hyperspectral images taken by two generations of X-ray telescopes with complementary spatial and spectral resolutions. Our method simultaneously deconvolves the instruments’ spatial and spectral responses, in order to return to the best resolutions of the two telescopes.
We implemented three regularisation schemes, the first being a term to promote sparsity under a 2D-1D wavelet transform, the second being a low rank approximation with Sobolev regularization (inspired by Guilloteau et al., 2020), and the third being a low rank approximation with a 2D wavelet sparsity term.
We tested our method on five toy-models, constructed from hydrodynamic simulations of supernova remnants by Orlando et al. (2016) convolved with an increasing complexity of instrumental response. Spatial features were well reconstructed, apart from a slight leakage in the bright shell in the toy-model with rebinning. The reconstruction of spectral features depended on the regularizations and model.
We found that while low rank methods accelerated the algorithm by half, they struggled to account for the strong spectral variability present around the iron line. The 2D-1D wavelet method accounted better for this variability, but it was somewhat biased in reconstructing the realistic model.
In conclusion, we believe that hyperspectral fusion shows promise and would greatly benefit the analysis of XRISM/Resolve data, though it would be useful to explore other regularization terms more suited to the variabilities of the data. In particular, we intend to combine fusion with a spectral physical model and source separation algorithm in future works. While the purpose of this article was a proof of concept, we are now actively working on applying this method to real data.

### v2

@@ 6 Conclusions
We have developed an algorithm for fusing hyperspectral images taken by two generations of X-ray telescopes with complementary spatial and spectral resolutions. Our method simultaneously deconvolves the instruments’ spatial and spectral responses in order to return to the best resolutions of the two telescopes.
We implemented three regularization schemes, the first being a term to promote sparsity under a 2D-1D wavelet transform, the second being a low-rank approximation with Sobolev regularization (inspired by Guilloteau et al., 2020), and the third being a low-rank approximation with a 2D wavelet sparsity term.
We tested our method on five toy models constructed from hydrodynamic simulations of SNRs by Orlando et al. (2016) and convolved with an increasing complexity of instrumental response. Spatial features are well reconstructed, apart from a slight leakage in the bright shell in the toy model with rebinning. The reconstruction of spectral features depends on the regularizations and model.
We find that while low-rank methods accelerated the algorithm by half, they struggled to account for the strong spectral variability present around the iron line. The 2D-1D wavelet method better accounts for this variability, but it is somewhat biased in reconstructing the realistic model.
In conclusion, we believe that hyperspectral fusion shows promise and would greatly benefit the analysis of XRISM/Resolve data; though it would be useful to explore other regularization terms more suited to the variabilities of the data. In particular, we intend to combine fusion with a spectral physical model and source-separation algorithm in future works. While the purpose of this article is to present a proof of concept, we are now actively working on applying this method to real data.
@@ Appendix A Metrics definition
Appendix B Additional Figures
This section presents additional figures for the results of section 4. Fig. 8 shows the obtained amplitude for for the three potential regularization, for each of the four toy models. For these same options, Fig. 9 shows the spectral angle maps. Fig. 10 displays the relative error maps averaged per pixel.
(a) Gaussian model between 0.5-1.4 keV, (b) Gaussian model between 6.2-6.9 keV, (c) Gaussian model with rebinning between 0.5-1.4 keV, and (d), Realistic model between 0.5-1.4 keV.
(a) Gaussian model between 0.5-1.4 keV, (b) Gaussian model between 6.2-6.9 keV, (c) Gaussian model with rebinning between 0.5-1.4 keV, and (d), Realistic model between 0.5-1.4 keV.
(a) Gaussian model between 0.5-1.4 keV, (b) Gaussian model between 6.2-6.9 keV, (c) Gaussian model with rebinning between 0.5-1.4 keV, and (d), Realistic model between 0.5-1.4 keV.